. Calculate the energy in eV of an electron in the 4th orbital of a helium atom
1
Expert's answer
2020-07-06T15:40:43-0400
"E=-13.6\frac{Z^2}{n^2}\\E=-13.6\frac{2^2}{4^2}=-3.4\ eV"